Peter was tasked to clean the library as his detention punishment. As he was cleaning the floor he pulled a vacuum cleaner with F= ( 40 i hat + 30 j hat) N and displaced at delta r = ( 3.0 i hat + 0.0 j hat) m.

A. Calculate the magnitudes of the displacement and the force
B. Calculate the work done by F
C. Find the angle between two vectors.
1
Expert's answer
2018-11-21T14:56:10-0500

F=√(〖40〗^2+〖30〗^2 )=50 N

δr=√(〖3.0〗^2+〖0.0〗^2 )=3.0 m

Work done

W=(F∙δr)=40×3.0+30×0.0=120 J

cos⁡〖θ=W/(F∙δr)〗=120/(50×3.0)=0.8

θ=37°